Select all of the linear equations in the list below.

A. y=π⋅x+2 B.5⋅y+2⋅x=7 C. y=5⋅x2+7 D. y2=3⋅x+4
E. y=3⋅x+2 F. y=2⋅x+7 G. y2=x+3 H. y=52⋅x I.2y=x+3
J. y=2⋅x+7 K. y+π=x-3

Answer by stanbon(75887)

Select all of the linear equations in the list below.
A. y=π⋅x+2 B. 5⋅y+2⋅x=7 C. y=5⋅x^2+7 D. y2=3⋅x+4
E. y=3⋅x+2 F. y=2⋅x+7 G. y^2=x+3 H. y=52⋅x I. 2y=x+3
J. y=2⋅x+7 K. y+π=x-3
----
Linears: A,B,E,F,H,I,J,K
-----
Cheers,
Stan H.